a. Choose a third animal. Calculate its speed in yards per second. Compare its speed with the speed of the other two animals you chose in Activity 1.
the first two animals i chose were :
fox 20.53 yards per second
and a coyote 21.02 yards per second
the third animal i chose is Zebra travels 40 m/hr (19.56 yd/sec)

a. Compare the zebra's speed with the speed of the two other animals
b. Decide which two animals should get head starts and how much of a head start each should get. Write three equations relating distance from the starting line to time.
c. Graph all three equations on the same coordinate plane.


I'm really stuck here and don't know what to do, could anyone please help me and also explain how i should do this? especially c. how should i graph them?

I would really appreciate it if you would help me out.

Answer by stanbon(75887)

a. Choose a third animal. Calculate its speed in yards per second. Compare its speed with the speed of the other two animals you chose in Activity 1.
the first two animals i chose were :
fox 20.53 yards per second
and a coyote 21.02 yards per second
the third animal i chose is Zebra travels 40 m/hr (19.56 yd/sec)
a. Compare the zebra's speed with the speed of the two other animals
zebra-coyote = 19.56-21.02 = -1.46 yd/sec
zebra-fox = 19.56 - 20.53 = -0.97 yd/
--------
b. Decide which two animals should get head starts and how much of a head start each should get.
If you want the race to be fair, the slowest animal needs the greater headstart.
Zebra is slowest: 19.56 yd/sec
then fox is next slowest: 20.53 yd/sec
coyote is fastest:21.02 yd/sec
-------
How much of a head start depends on the length of the race
If the race is 100 yds, coyote will take dist/rate = 100/21.02 = 4.57 seconds
fox will take 100/20.53 = 4.87 sec
----
So the fox needs a 4.87-4.57 = 0.3 sec head start.
At the fox's rate that is a head start of 0.3*20.53 = 6.159 yards
----
Do the same for the zebra to figure out it's head start.
=======
Write three equations relating distance from the starting line to time.
c. Graph all three equations on the same coordinate plane.
---
Cheers,
Stan H.
===============